AVZ is a​ start-up company who is using all its cash to growth so it does not plan to pay dividends for the next 6 years. The company then plans to start paying annual cash dividends starting in year 7 of ​$6.00 for 10 years. ​ Thereafter, the company will assume a constant growth dividend policy and the estimated growth rate in dividends forever after that point is 2​%. The price of the stock is set to yield a return of 12​%. What is the price of this stock​ today?
Answer #1

PV of annuity at time 6 = Annuity*(1-1/(1+rate)^number of terms)/rate

= 6*(1-1/1.12^10)/0.12

=33.90

Terminal value at year 16 = D17/(k-g)

= 6*102%/(12%-2%)

= 61.2

Current stock price = PV of dividends + PV of terminal value

= 33.90/1.12^6+61.2/1.12^16

=$27.16